The rotfloat package^{∗}

Harald Axel Sommerfeldt axel@hp1.ang-physik.uni-kiel.de

1995/03/30

Abstract

The float package [1] provides commands to define new floats of various styles (plain, boxed, ruled, and userdefined ones ); the rotating package [2] provides new environments (sidewaysfigure and sidewaystable) which are rotated by 90^{∘} or 270^{∘}. But what about new rotated floats, e.g. a rotated ruled one? This package makes this possible; it builds a bridge between both packages and extend the commands from the float package to define rotated versions of the new floats, too.

1 The user interface

To use this package just type

     \usepackage[⟨options⟩]{rotfloat}

in the preamble of your document. The options are exactly the same as for the rotating package, in fact this package hasn’t got any own options at all.

\newfloat The commands \newfloat and \restylefloat from the float package (re)define \restylefloat the float type ⟨type⟩ and now additionally a rotated one called ⟨sidewaystype⟩:

\newrotfloat{⟨type⟩}{⟨placement⟩}{⟨ext⟩}[⟨within⟩]

     \restylerotfloat{⟨type⟩}

E.g. the code

     \floatstyle{ruled}
     \floatname{program}{Program}
     \newfloat{program}{tbp}{lop}[section]

defines the new floating environments program, program*, and sidewaysprogram which behave equivalent to figure, figure*, and sidewaysfigure.

∗This package has version number 1.0, last revised 1995/03/30.

The code

     \floatstyle{boxed}
     \restylefloat{table}

will restyle the environments table, table*, and sidewaystable. Please take a look at the float package for a complete description of these commands. Additionally an example file is provided with this package.

2 A final note

This package was written and tested with the version 1.2 of the float package and version 2.6 of the rotating package. You should not use this package together with older versions of them! Future versions of these packages may make some trouble, but I hope they will not.

References

[1] Anselm Lingnau: An Improved Environment for Floats, 1994/06/20

[2] Sebastian Rahtz and Leonor Barroca: A style option for rotated objects in L^{A}TEX, 1994/08/26

Change History

v1.0 General: First release . . . . . . . . . . 1

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

N

\newfloat . . . . . . . . . 1

R

                          \restylefloat . . . . . . 1